OpenStudy (anonymous):

What is the partial fraction expansion of (xsqrd-8x)/[(x+1)(x+4)^3]?

OpenStudy (anonymous):

is it \[ \frac { x^2 -8x}{ (x-1)(x+4)^3} \]

OpenStudy (anonymous):

Yeah that's the function. I'm trying to find the expansion so I can take the indefinite integral.

OpenStudy (anonymous):

Oh okay thanks, It looks like A= 1/3 and B=-16 via cover up method. But I could be wrong. I'm not sure how to find the other values though. Any advice?

OpenStudy (anonymous):

No A is not 1/3. Try again.

OpenStudy (anonymous):

Idk I keep getting 1/3 :/

OpenStudy (anonymous):

What happened to the x in front of the 8? On the top polynomial of the rational function?

OpenStudy (anonymous):